Use the tables to answer questions. One student from each school is chosen at random to be on a committee. Find each probability.
School A
Freshman Sophomores Juniors Seniors
30% 27% 25% 18%
School B
Freshman Sophomores Juniors Seniors
28% 28% 24% 20%
1. junior from School A and a senior from School B
2. Two junior
3. A freshman or sophomore from School A and a senior from School B

Answer by stanbon(57361)

Use the tables to answer questions. One student from each school is chosen at random to be on a committee. Find each probability.
School A
Freshman Sophomores Juniors Seniors
30% 27% 25% 18%
School B
Freshman Sophomores Juniors Seniors
28% 28% 24% 20%
1. P(junior from School A and a senior from School B) = 0.25*0.20 = 0.05
------
2. P(Two juniors) = 0.25*0.24 = 0.06
-----------
3. P(A freshman or sophomore from School A and a senior from School B)
= 0.57*0.20 = 0.114
-----------------------------
Cheers,
Stan H.
===========